OpenStudy (anonymous):

I'm studying for my final tomorrow and I need to see how you would solve... Log10^(x+4) + log10^(x-2)=1

OpenStudy (anonymous):

This can be rewritten: \[\log(x) + \log(y) = \log(xy)\]

OpenStudy (anonymous):

And why is that?

OpenStudy (anonymous):

The long and short of it is: because thats how logs work. Without some long formal proof.

OpenStudy (anonymous):

Ahhh.. Got it.. Thank you!

OpenStudy (anonymous):

No prob.